# The radiation condition for Helmholtz equations above locally perturbed periodic surfaces

### v1

@@ 9 Radiation condition for $u_{2}$
We have considered $v_{1}$ defined by (4.7) where $u$ is given by (4.5) in the previous section. In this section, we will consider $v_{2}$ be defined by (4.7) where $u$ be given by (4.6). Following the arguments at the beginning of Section 8, we only need to consider the radiation conditions for functions defined
@@
$\psi(\theta)=-k^{3/2}\sin\theta\phi(k\cos\theta-j)\left|\frac{\cos\theta-\cos\theta_{0}}{\theta-\theta_{0}}\right|\in C_{0}^{\infty}(\beta_{1},\beta_{2}).$
@@
Use the formula $\cos(\theta-\theta_{*})=1-2\sin^{2}((\theta-\theta_{*})/2)$ and let $t:=\sin((\theta-\theta_{*})/2)$, then $\theta=2\arcsin t+\theta_{*}$ and
@@
$\zeta(t)$ $=\frac{2}{\sqrt{1-t^{2}}}\left|\frac{2\arcsin t+\theta_{*}-\theta_{0}}{t-a}\right|\psi(2\arcsin t+\theta_{*})$
$=\frac{2}{\sqrt{1-t^{2}}}\frac{2\arcsin t+\theta_{*}-\theta_{0}}{t-a}\psi(2\arcsin t+\theta_{*})\in C_{0}^{\infty}(\gamma_{1},\gamma_{2}).$
Since the formula has exactly the same form as $I^{D}_{1}(r)$ defined by (5.18), from Theorem 5.9 we have
@@
Following Section 8.1, with integration by parts,
@@
$:=\frac{1}{r}\int_{\beta_{1}}^{\beta_{2}}|\theta-\theta_{0}|\zeta_{1}(\theta)e^{\mathrm{i}kr\cos(\theta-\theta_{*})}\,\mathrm{d}\theta\,+\frac{1}{r}\int_{\beta_{1}}^{\beta_{2}}{\rm sgn}(x-x_{0})\zeta_{1}(\theta)e^{\mathrm{i}kr\cos(\theta-\theta_{*})}\,\mathrm{d}\theta\,$
@@
and both functions belong to $C_{0}^{\infty}(\gamma_{1},\gamma_{2})$. Since $(I)$ has the form of $I^{D}_{1}(r)$ defined by (5.18) and $(II)$ has the form of $I^{D}_{2}(r)$ define by (5.20), from the results in Theorem 5.9 and Corollary 5.10,
@@
holds uniformly for all $j\in J_{+}(\alpha_{0})$.
@@
$\psi(\theta)=2k^{3/2}\phi(k\cos\theta-j)\frac{\sin^{2}(\theta/2)}{\sin^{2}\theta}\in C^{\infty}[0,\beta)$
@@
$\frac{\partial v}{\partial r}-\mathrm{i}kv=\mathrm{i}k\int_{0}^{\beta}\sin^{3}\theta\left(\cos(\theta-\theta_{*})-1\right)\psi(\theta)e^{\mathrm{i}kr\cos(\theta-\theta_{*})}\,\mathrm{d}\theta\,.$ (9.9)
@@
$|v(x)|\leq Cr^{-1/2},\quad\left|\frac{\partial v}{\partial r}-\mathrm{i}kv\right|\leq Cr^{-3/2}.$ (9.10)
@@
$|v(x)|,\,\left|\frac{\partial v}{\partial r}-\mathrm{i}kv\right|\leq Cr^{-2}.$ (9.13)
@@ 10 Proof for Part (B) of the Main Theorem
Now we conclude the proof for Part (B) of the main theorem, i.e., Theorem 4.2. The results are concluded from Section 7, 8 and 9. According to Section 3.2, the structures of the solutions are different depending on the wavenumber $k$, thus we will prove the two cases separately.
@@
According to (3.7), we apply the Floquet-Bloch transform for function $u\big{|}_{\Gamma^{H}}$ and the transformed field has the form of
$w_{0}(\alpha,x)+\sqrt{\alpha-\alpha_{1}}w_{1}(\alpha,x)+\sqrt{\alpha-\alpha_{2}}w_{2}(\alpha,x)$
where $w_{0}\in C^{\infty}_{\mathrm{per}}\left(\Lambda;H^{1}_{\alpha}(\Gamma^{H}_{0})\right)$, and $w_{1},w_{2}\in C_{0}^{\infty}\left((\alpha_{j}-\delta,\alpha_{j}+\delta);H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Then we get back to $u\big{|}_{\Gamma^{H}}$ by the inverse Floquet-Bloch transform:
@@
When $j\in J_{-}^{1}$, with Theorem 8.1, 8.2 and 8.3, there is a constant $C>0$ such that
@@
holds uniformly for $x\in\mathbb{R}^{+}_{H+h}$.
When $j\in J_{0}^{1}$, with Theorem 8.6 and 8.7, there is a constant $C>0$ such that
@@
holds uniformly for $x\in\mathbb{R}^{+}_{H+h}$.
When $j\in J_{+}^{1}$, with Theorem 8.4 and 8.5, there are two constants $C,{\varepsilon}_{0}>0$ such that
@@
According to (3.8), the Floquet-Bloch transform of $u\big{|}_{\Gamma^{H}}$ has the form of
$w_{0}(\alpha,x)+\sqrt{\alpha-\alpha_{0}}w_{1}(\alpha,x)+|\alpha-\alpha_{0}|w_{2}(\alpha,x)$
where $w_{0}\in C^{\infty}_{\mathrm{per}}\left(\Lambda;H^{1}_{\alpha}(\Gamma^{H}_{0})\right)$, and $w_{1},w_{2}\in C_{0}^{\infty}\left((\alpha_{j}-\delta,\alpha_{j}+\delta);H^{1}_{\alpha}(\Omega^{H}_{0})\right)$. Then we get back to $u\big{|}_{\Gamma^{H}}$ by the inverse Floquet-Bloch transform:
@@
When $j\in J_{-}$, with Theorem 9.1 and 9.2, there is a constant $C>0$ such that
@@
holds uniformly for $x\in\mathbb{R}^{+}_{H+h}$.
When $j\in J_{0}$, with Corollary 9.5, 9.6 and 9.7, there is a constant $C>0$ such that
@@
holds uniformly for $x\in\mathbb{R}^{+}_{H+h}$.
When $j\in J_{+}^{1}$, with Theorem 9.3 and 9.4, there are two constants $C,{\varepsilon}_{0}>0$ such that
@@ 11 Radiation condition for scattering problems with locally perturbed periodic surfaces
In this section, we extend the results to the case when the periodic surface is perturbed. Let $p:\,\mathbb{R}\rightarrow\mathbb{R}$ be a function which is compactly supported. Assume that $p=0$ when $|x_{1}|\geq L$ for some $L>0$. Let
@@
Moreover, $u_{T}$ satisfies the problem
@@
where $u_{r}$ satisfies the Rayleigh expansion (3.6) and $u_{d}$ satisfies the half-space
Appendix
Fresnel integral and its application
Recall the Fresnel integrals (see [1])
$S(t)=\int_{0}^{t}\sin(z^{2})\,\mathrm{d}z\,,\quad C(t)=\int_{0}^{t}\cos(z^{2})\,\mathrm{d}z\,$ (11.7)
@@
$\int_{0}^{\infty}\sqrt{x}e^{\mathrm{i}x^{2}}\,\mathrm{d}x\,=\frac{1}{2}\Gamma\left(\frac{3}{4}\right)\exp\left(\frac{3\mathrm{i}\pi}{8}\right).$ (11.8)
Properties of the Bessel functions
@@
$\left(\frac{1}{z}\frac{\,\mathrm{d}missing\,}{\,\mathrm{d}z\,}\right)^{m}[z^{n}J_{n}(z)]=z^{n-m}J_{n-m}(z).$ (11.9)
@@
$\frac{\,\mathrm{d}missing\,}{\,\mathrm{d}x\,}[x^{n}J_{n}(rx)]=rx^{n}J_{n-1}(rx).$ (11.10)
@@
$\frac{\,\mathrm{d}missing\,}{\,\mathrm{d}t\,}H_{0}^{(1)}(t)=-H_{1}^{(1)}(t);$ (11.11)
$\frac{\,\mathrm{d}missing\,}{\,\mathrm{d}t\,}H_{1}^{(1)}(t)=H_{0}^{(1)}(t)-\frac{1}{t}H_{1}^{(1)}(t).$ (11.12)
Second property is the envelope of the Bessel functions. From [23], it is estimated that
$|J_{n}(x)|\leq c_{0}|x|^{-1/3},$ (11.13)
where $c_{0}=0.7857468704\dots$.
We have more accurate error estimations for Bessel/Hankel functions, see [31, 3].
First for any natural number $n$, define the polynomial $a_{k}(n)$ as:
$a_{0}(n)=1,\quad a_{k}(n)=\frac{(4n^{2}-1)(4n^{2}-3)\cdots(4n^{2}-(2k-1)^{2})}{k!8^{k}}.$
The Hankel functions of the first and second kind has the form of
$H_{n}^{(1)}(x)=\sqrt{\frac{2}{\pi x}}\exp\left(\mathrm{i}\left(x-\frac{n\pi}{2}-\frac{\pi}{4}\right)\right)\left[\sum_{j=0}^{\ell-1}\mathrm{i}^{j}\frac{a_{j}(n)}{x^{j}}+R_{\ell}^{+}(n,x)\right];$ (11.14)
$H_{n}^{(2)}(x)=\sqrt{\frac{2}{\pi x}}\exp\left(-\mathrm{i}\left(x-\frac{n\pi}{2}-\frac{\pi}{4}\right)\right)\left[\sum_{j=0}^{\ell-1}(-\mathrm{i})^{j}\frac{a_{j}(n)}{x^{j}}+R_{\ell}^{-}(n,x)\right];$ (11.15)
where $x>0$ takes the positive value. Here $R_{\ell}^{\pm}(n,x)$ are estimated by:
$\left|R_{\ell}^{\pm}(n,x)\right|\leq 2|a_{\ell}(n)|x^{-\ell}\exp\left(|n^{2}-1/4|x^{-1}\right).$
In particular, take $\ell=1$, then we have:
$H_{n}^{(1)}(x)=\sqrt{\frac{2}{\pi x}}\exp\left(\mathrm{i}\left(x-\frac{n\pi}{2}-\frac{\pi}{4}\right)\right)\left[1+R_{1}^{+}(n,x)\right];$
$H_{n}^{(2)}(x)=\sqrt{\frac{2}{\pi x}}\exp\left(-\mathrm{i}\left(x-\frac{n\pi}{2}-\frac{\pi}{4}\right)\right)\left[1+R_{\ell}^{-}(n,x)\right];$
@@
$\left|R_{1}^{\pm}(n,x)\right|$ $\leq 2|a_{\ell}(n)||x|^{-1}\exp\left(|n^{2}-1/4||x|^{-1}\right)$
$\leq|n^{2}-1/4|x^{-1}\exp\left(|n^{2}-1/4|x^{-1}\right)$
and $C>0$ is a constant Here we used the fact that
$a_{1}(n)=\frac{4n^{2}-1}{8}=\frac{1}{2}(n^{2}-1/4).$
Thus
$J_{n}(x)$ $=\frac{H_{n}^{(1)}(x)+H_{n}^{(2)}(x)}{2}$
$=\sqrt{\frac{2}{\pi x}}\cos\left(x-\frac{n\pi}{2}-\frac{\pi}{4}\right)\left[1+R(n,x)\right],$ (11.16)

### v2

@@ 8 Radiation condition for $u_{2}$
In this section, we will consider $v_{2}$ be defined by (4.6) where $u$ be given by (4.5). Following the arguments at the beginning of Section 7, we only need to consider the radiation conditions for functions defined
@@
$\psi(\theta)=-k^{2}\sin\theta\phi(k\cos\theta-j)\left|\frac{\cos\theta-\cos\theta_{0}}{\theta-\theta_{0}}\right|\in C_{0}^{\infty}(\beta_{1},\beta_{2}).$
@@
Use the formula $\cos(\theta-\theta_{*})=1-2\sin^{2}((\theta-\theta_{*})/2)$ and let $t:=\sin((\theta-\theta_{*})/2)\in[\gamma_{1},\gamma_{2}]\subset(-1,1)$, then $\theta=2\arcsin t+\theta_{*}$ and
@@
$\zeta(t)$ $=\frac{2}{\sqrt{1-t^{2}}}\left|\frac{2\arcsin t-2\arcsin a}{t-a}\right|\psi(2\arcsin t+\theta_{*})$
$=\frac{2}{\sqrt{1-t^{2}}}\frac{2\arcsin t-2\arcsin 2}{t-a}\psi(2\arcsin t+\theta_{*})\in C_{0}^{\infty}(\gamma_{1},\gamma_{2}).$
Since the formula has exactly the same form as $I^{D}_{1}(r)$ defined by (B.18), from Theorem B.9 we have
@@
$\left|\frac{\partial v}{\partial r}-\mathrm{i}kv\right|\leq Cr^{-3/2}$ (8.5)
holds uniformly for $x\in U_{H+h}$.
@@
Following Theorem 7.2, with integration by parts,
@@
$:=\frac{1}{r}\int_{\beta_{1}}^{\beta_{2}}|\theta-\theta_{0}|\zeta_{1}(\theta)e^{\mathrm{i}kr\cos(\theta-\theta_{*})}\,\mathrm{d}\theta\,+\frac{1}{r}\int_{\beta_{1}}^{\beta_{2}}{\rm sgn}(\theta-\theta_{0})\zeta_{2}(\theta)e^{\mathrm{i}kr\cos(\theta-\theta_{*})}\,\mathrm{d}\theta\,$
@@
and both functions belong to $C_{0}^{\infty}(\gamma_{1},\gamma_{2})$.
For both integrals, let $t:=\sin((\theta-\theta_{*})/2)$ and $a:=\sin((\theta_{0}-\theta_{*})/2)$, then
$(I)$ $=e^{\mathrm{i}rt}\int_{\gamma_{1}}^{\gamma_{2}}|t-a|\rho_{1}(t)e^{-2\mathrm{i}krt^{2}}\,\mathrm{d}t\,;$
$(II)$ $=e^{\mathrm{i}rt}\int_{\gamma_{1}}^{\gamma_{2}}{\rm sgn}(t-a)\rho_{2}(t)e^{-2\mathrm{i}krt^{2}}\,\mathrm{d}t\,;$
where
$\rho_{1}(t)$ $=\frac{2}{\sqrt{1-t^{2}}}\frac{2\arcsin t-2\arcsin a}{t-a}\zeta_{1}(2\arcsin t+\theta_{*})\in C_{0}^{\infty}(\gamma_{1},\gamma_{2});$
$\rho_{2}(t)$ $=\frac{2}{\sqrt{1-t^{2}}}\zeta_{2}(2\arcsin t+\theta_{*})\in C_{0}^{\infty}(\gamma_{1},\gamma_{2}).$
Since $(I)$ has the form of $I^{D}_{1}(r)$ defined by (B.18) and $(II)$ has the form of $I^{D}_{2}(r)$ define by (B.20), from the results in Theorem B.9 and Corollary B.10,
@@
holds uniformly for all $j\in J_{+}(\alpha_{0})$ and $x\in U_{H+h}$.
@@
$\psi(\theta)=2k^{2}\phi(k\cos\theta-j)\frac{\sin^{2}(\theta/2)}{\sin^{2}\theta}\in C^{\infty}[0,\beta)$
@@
$\frac{\partial v}{\partial r}-\mathrm{i}kv=\mathrm{i}k\int_{0}^{\beta}\sin^{2}\theta\sin\theta\left(\cos(\theta-\theta_{*})-1\right)\psi(\theta)e^{\mathrm{i}kr\cos(\theta-\theta_{*})}\,\mathrm{d}\theta\,.$ (8.9)
@@
$|v(x)|\leq Cr^{-1/2},\quad\left|\frac{\partial v}{\partial r}-\mathrm{i}kv\right|\leq Cr^{-3/2}$ (8.10)
holds uniformly for $x\in U_{H+h}$
@@
$|v(x)|,\,\left|\frac{\partial v}{\partial r}-\mathrm{i}kv\right|\leq Cr^{-2}$ (8.13)
holds uniformly for $x\in U_{H+h}$.
@@
$|v(x)|,\,\left|\frac{\partial v}{\partial r}-\mathrm{i}kv\right|\leq Cr^{-3/2}$ (8.14)
holds uniformly for $x\in U_{H+h}$.
@@ 9 Proof for Part (B) of the Main Theorem
Now we conclude the proof for Part (B) of the main theorem, i.e., Theorem 4.2. The results are concluded from Section 6, 7 and 8. According to Theorem 3.5, the structures of the solutions are different depending on the wavenumber $k$, thus we will prove the two cases separately.
@@
We recall that from the inverse Floquet-Bloch transform and the regularity of $w$ given by (3.7), $u$ is given by (3.8)
@@
When $j\in J_{-}^{1}$, with Theorem 7.1, 7.2 and 7.3, there is a constant $C>0$ such that
@@
holds uniformly for $x\in U_{H+h}$.
When $j\in J_{0}^{1}$, with Theorem 7.6 and 7.7, there is a constant $C>0$ such that
@@
holds uniformly for $x\in U_{H+h}$.
When $j\in J_{+}^{1}$, with Theorem 7.4 and 7.5, there are two constants $C,{\varepsilon}_{0}>0$ such that
@@
We recall that from the inverse Floquet-Bloch transform and the regularity of $w$ given by (3.9), $u$ is given by (3.10)
@@
When $j\in J_{-}$, with Theorem 8.1 and 8.2, there is a constant $C>0$ such that
@@
holds uniformly for $x\in U_{H+h}$.
When $j\in J_{0}$, with Corollary 8.5, 8.6 and 8.7, there is a constant $C>0$ such that
@@
holds uniformly for $x\in U_{H+h}$.
When $j\in J_{+}^{1}$, with Theorem 8.3 and 8.4, there are two constants $C,{\varepsilon}_{0}>0$ such that
@@ 10 Radiation condition for scattering problems with locally perturbed periodic surfaces
In this section, we extend the results to the case when the periodic surface is perturbed. Let $p:\,\mathbb{R}\rightarrow\mathbb{R}$ be a function which is compactly supported, then there is an $L>0$ such that $p(x_{1})=0$ when $|x_{1}|\geq L$. Let
@@
Thus the radiation condition for $u_{T}$ is exactly the same as $u_{T}$. Moreover, $u_{T}$ satisfies the problem
@@
where $u_{r}$ satisfies the Rayleigh expansion (3.6) and $u_{d}$ satisfies the radiation condition defined in Definition 4.1. The result is now summarized as the following theorem.
Theorem 10.1.
Let $\Gamma_{p}$ be a local perturbation of the periodic surface $\Gamma$. Given the incident field $u^{i}$, which is either the point source, i.e., $u^{i}=G(x,y)$, or the plane wave $u^{i}=e^{\mathrm{i}k(x_{1}\cos\theta-x_{2}\sin\theta)}$ with $\theta\in(0,\pi)$. Then
  1. (i)
    when $u^{i}$ is a point source, the scattered field $u_{p}$ satisfies the radiation condition defined by Definition 4.1;
  2. (ii)
    when $u^{i}$ is a plane wave, the scattered field $u_{p}$ is decomposed as
    $u_{p}=u_{r}+u_{d},$
    where $u_{r}$ satisfies the Rayleigh expansion (3.6) with $\alpha=k\cos\theta$ and $u_{d}$ satisfies the radiation condition defined by Definition 4.1.
Appendix A Appendix A: Useful formulas and estimations
A.1 Fresnel integral and its application
Recall that the Fresnel integrals are defined by (see [1]):
$S(t)=\int_{0}^{t}\sin(z^{2})\,\mathrm{d}z\,,\quad C(t)=\int_{0}^{t}\cos(z^{2})\,\mathrm{d}z\,$ (A.1)
@@ A.1 Fresnel integral and its application
$\int_{0}^{\infty}\sqrt{x}e^{\mathrm{i}x^{2}}\,\mathrm{d}x\,=\frac{1}{2}\Gamma\left(\frac{3}{4}\right)\exp\left(\frac{3\mathrm{i}\pi}{8}\right).$ (A.2)
A.2 Properties of the Bessel functions
@@
$\left(\frac{1}{z}\frac{\,\mathrm{d}missing\,}{\,\mathrm{d}z\,}\right)^{m}[z^{n}J_{n}(z)]=z^{n-m}J_{n-m}(z).$ (A.3)
@@
$\frac{\,\mathrm{d}missing\,}{\,\mathrm{d}x\,}[x^{n}J_{n}(rx)]=rx^{n}J_{n-1}(rx).$ (A.4)
@@
$\frac{\,\mathrm{d}missing\,}{\,\mathrm{d}t\,}H_{0}^{(1)}(t)=-H_{1}^{(1)}(t);$ (A.5)
$\frac{\,\mathrm{d}missing\,}{\,\mathrm{d}t\,}H_{1}^{(1)}(t)=H_{0}^{(1)}(t)-\frac{1}{t}H_{1}^{(1)}(t).$ (A.6)
We also need the asymptotic behaviour
$H_{n}^{(1)}(t)=\sqrt{\frac{2}{\pi t}}\exp\left(\mathrm{i}\left(t-\frac{n\pi}{2}-\frac{\pi}{4}\right)\right)\left[1+O\left(\frac{1}{t}\right)\right],\quad t\rightarrow\infty.$ (A.7)
Appendix B Appendix B: Analysis of some important integrals
To prepare for the estimation of the radiation condition in Theorem 4.2 with the help of the Floquet-Bloch transform, we need to estimate the decay of some integrals depends on parameters, which are concluded in Table B.
class name definition decay rate
We will consider the decay rate of the above integrals when $r\rightarrow\infty$. Here $\delta>0$ is a fixed small value. For simplicity, we all require that $\phi\in C^{\infty}_{0}(\mathbb{R})$ and vanishes either near $\delta$ (for example in $I^{A}_{1}(r),\,I^{A}_{2}(r),\,I^{B}_{1}(r),\,I^{B}_{2}(r)$) or the edge points $\gamma_{1},\gamma_{2}$ (for example in $I^{C}_{1}(r),\,I^{C}_{2}(r),\,I^{D}_{1}(r),\,I^{D}_{2}(r)$. Note that the results can be naturally extended to more general functions, i.e., the condition $\phi\in C^{\infty}_{0}(\mathbb{R})$ is always too strong (for example see Remark B.2). We will not discuss the very explicit explanations since it will be too long for the readers. For the discussions of the integrals of Classes A, B, C and D, we refer to Subsections B.1, B.2, B.3 and B.4, respectively.
B.1 Analyse of the integrals of Class A
We begin with the first integral of Class A:
$I_{1}^{A}(r):=\int_{0}^{\delta}\frac{1}{\sqrt{x}}\phi(x)e^{\mathrm{i}rx}\,\mathrm{d}x\,,\quad r>0.$ (B.1)
Theorem B.1.
$|I^{A}_{1}(r)|\leq Cr^{-1/2},\quad\forall\,r>0.$ (B.2)
Proof.
From direct computation,
@@ B.1 Analyse of the integrals of Class A
where $S(x)$ and $C(t)$ are Fresnel integrals (see equation A.1) and uniformly bounded. Thus
For (II), with integration by parts,